While skydiving you let 35s pass before opening your parachute. How fast are you traveling
1
Expert's answer
2019-09-17T12:07:57-0400

The velocity of a skydiver after 35 s


"v=v_i+gt=0+9.8\:\rm m/s^2\times 35\: s=343\: m/s"